R=1Kohm resistor ,a C =1uF capacitor and L =0.2H inductorare connected in series with V=150sin(377t) volt source what is the maximum current delievered
1
Expert's answer
2020-11-06T10:12:33-0500

Maximum current for serial RLC circuit:

"I=\frac{U}{\sqrt{R^2+(\omega\times L-\frac{1}{\omega\times C})^2}}"

where U - maximum voltage, R - passive resistance, "\omega" - angular frequency, L - inductivity, C - capacity

"I=\frac{150}{\sqrt{1000^2+(377\times 0.2-\frac{1}{377\times 1\times10^{-6}})^2}}"

"I=\frac{150}{\sqrt{10^6+(75.4-2652.5)^2}}=\frac{150}{\sqrt{10^6+6641444.41}}=\frac{150}{2764}=0.054A"